OpenStudy (anonymous):

log{1/3(a+b)} = 1/2(log a + log b) then a^2 + b^2 = ?

OpenStudy (anonymous):

any1 answer

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

u want options : a) 9ab b) 7ab c) 5ab d) none of these

OpenStudy (anonymous):

loga+logb=logab 1/2log(ab)=logab^1/2 now (ab)^1/2=1/3(a+b) a+b=3(ab)^1/2 (a+b)^2=9ab a^2+b^2=(a+b)^2 - 2ab hence 9ab -2ab =7ab

OpenStudy (anonymous):

did you get it

OpenStudy (nilankshi):

good

OpenStudy (anonymous):

yes vicky thanks very much for ur support

OpenStudy (anonymous):

yes vicky thanks very much for ur support

OpenStudy (anonymous):

yes vicky thanks very much for ur support